OpenStudy (loser66):

Use epsilon-delta to prove \[lim_{x\rightarrow -1}\dfrac{x+5}{2x+3}=4\] Please, help

OpenStudy (anonymous):

\[ |x-1|<\delta \implies \left|\frac{x+5}{2x+3}-4\right|<\epsilon \]

OpenStudy (anonymous):

You start here.

OpenStudy (loser66):

no, |x+1|, not x-1

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

\[ |x-(-1)|<\delta \implies \left|\frac{x+5}{2x+3}-4\right|<\epsilon \]

OpenStudy (loser66):

yes, but I got stuck at argue \(|\dfrac{-7}{2x+3}| \)

OpenStudy (anonymous):

Can you simplify the inequalities a bit?

OpenStudy (loser66):

yes, it is \(|\dfrac{-7}{2x+3}||x+1|<\varepsilon\)

OpenStudy (anonymous):

Okay now then, You want to make sure that \[ \left|\frac{-7}{2x+3}\right|\delta <\epsilon \]Right?

OpenStudy (anonymous):

First thing you want to do is figure out what values the function will be when \(x\) is near \(-1\).

OpenStudy (loser66):

it is 4

OpenStudy (anonymous):

Alternatively, if there is a maximum value, you want to find the maximum value of it.

OpenStudy (loser66):

\(|x+1|<\delta \\-\delta< x+1<\delta\\-1-\delta<x<-1+\delta\) If \(\delta =1,\) then -2< x< 0 hence -4< 2x < 0 and -1< 2x +3 <3 That gives me \(\dfrac{1}{2x+3}< -1\) or \(\dfrac{1}{2x+3}>3\)

OpenStudy (loser66):

and how to put -7 in to those intervals, and take absolute value to get the expression < epsilon?

OpenStudy (anonymous):

Actually, you should look at this first: http://www.wolframalpha.com/input/?i=%7C-7%2F(2x%2B3)%7C

OpenStudy (anonymous):

I think a good minimum range for delta would be something like \(1/4\). If you get to \(-1.5\) then we get into trouble.

OpenStudy (anonymous):

So in the range: \[ -1/4 <x+1<1/4 \]What sort of range do we see for \[ \left|-\frac{7}{2x+3}\right| \]

OpenStudy (anonymous):

When \(\delta = 1/4\), when we see that max value of it will be \(14\).

OpenStudy (loser66):

Is there any logic way to have delta or just guess? Since we are not allowed to use math tool in class, how can I get that value?

OpenStudy (anonymous):

We can say \(\delta = \min(1/4, f(\epsilon))\) so that \(\delta\) neger gets to large.

OpenStudy (loser66):

What if 1/4 is too small?

OpenStudy (anonymous):

Now, if \(x\) gets really close to \(-1\), we know that \[ \left|\frac{-7}{2x+3}\right|<14 \]

OpenStudy (anonymous):

We have this relationship: \[ \delta <\frac{\epsilon }{\left|\frac{-7}{2x+3}\right|} \]

OpenStudy (anonymous):

\[ \delta <\frac{\epsilon}{14}\leq \frac{\epsilon }{\left|\frac{-7}{2x+3}\right|} \]

OpenStudy (anonymous):

So we can safely say \[ \delta = \min (1/4, \epsilon/14) \]

OpenStudy (loser66):

then, take delta = min {1/4, epsilon/14} and go backward, right?

OpenStudy (loser66):

I got it. Thanks a lot. :)

OpenStudy (anonymous):

I suppose so. That would be a big rigorous.

OpenStudy (loser66):

lost connection again :(